OpenStudy (anonymous):

What is the slope-intercept form of this equation? 6x + 4y = 8 What is the slope-intercept form of this equation? 6x + 4y = 8 What is the slope-intercept form of this equation? 6x + 4y = 8 What is the slope-intercept form of this equation? 6x + 4y = 8 A. y = -3/2x + 2 B. y = -3/2x + 4 C. y = 3/2x + 4 D. y = 3/4x + 2

OpenStudy (anonymous):

\[ 6x + 4y = 8 \implies 3x + 2y = 4 \implies 2y = 4 - 3x \implies y = 2 - \frac {3}{2} x, \]so the answer is \[ \boxed{\textbf{(A)}} \]

OpenStudy (dmndlife24):